# Bag of Bags: Adaptive Visual Vocabularies for Genizah Join Image Retrieval

### 09 Apr 2026 v1

@@ Abstract
A join is a set of manuscript fragments identified as originally emanating from the same manuscript. We study manuscript join retrieval: Given a query image of a fragment, retrieve other fragments originating from the same physical manuscript. We propose Bag of Bags (BoB), an image-level representation that replaces the global-level visual codebook of classical Bag of Words (BoW) with a fragment-specific vocabulary of local visual words. Our pipeline trains a sparse convolutional autoencoder on binarized fragment patches, encodes connected components from each page, clusters the resulting embeddings with per image $k$-means, and compares images using set to set distances between their local vocabularies. Evaluated on fragments from the Cairo Genizah, the best BoB variant (viz. Chamfer) achieves Hit@1 of 0.78 and MRR of 0.84, compared to 0.74 and 0.80, respectively, for the strongest BoW baseline (BoW-RawPatches-$\chi^{2}$), a 6.1% relative improvement in top-1 accuracy. We furthermore study a mass-weighted BoB-OT variant that incorporates cluster population into prototype matching and present a formal approximation guarantee bounding its deviation from full component-level optimal transport. A two-stage pipeline using a BoW shortlist followed by BoB-OT reranking provides a practical compromise between retrieval strength and computational cost, supporting applicability to larger manuscript collections. The code and dataset are available at https://github.com/TAU-CH/midrash_bob.
@@ 1 Introduction
Automating join retrieval is a challenging computer vision task because many manuscript fragments are often severely damaged, incomplete, and stained. The difficulty arises from the fact that fragments originating from the same manuscript may share only subtle, handwriting style specific visual patterns, while fragments from different manuscripts can still appear globally similar in aspects such as background texture, aging, and degradation level.
Standard retrieval architectures frequently rely on Bag of Words (BoW) paradigms. BoW maps images to histograms over a shared global codebook. It summarizes all fragments with respect to the same dictionary. However, this global quantization is fundamentally limited for paleographic analysis, as it causes critical image specific handwriting style information to be lost during quantization. Two images with identical codeword frequencies but different geometric character style distributions are assigned distance zero by BoW, yet they may have been written by entirely different hands. Conversely, fragments from the same manuscript may differ in lighting and layout, causing BoW to miss them. BoW’s pooling over a global codebook erases the image specific handwriting style signal.
@@
  • We propose Bag of Bags (BoB), a fragment specific representation for fragment retrieval that replaces a shared global visual-word codebook with per-fragment vocabularies of local visual words, constructed from sparse autoencoder embeddings of connected-component patches.
@@ 2 Related Work
Automated identification of joins was first explored in [21] who proposed a method for finding candidate matches between manuscript fragments. Their approach represents each image using local visual features and a bag of features representation, and then compares pairs of leaves using a learned similarity measure to determine whether they may belong to the same original work. However, this method requires computationally expensive pairwise comparisons during retrieval. Analogous fragment assembly problems arise in other historical document collections, including ancient papyri, where AI-based approaches have been surveyed for matching dispersed fragments via visual texture and learned similarity measures [20, 16].
@@
More broadly, historical document retrieval has been studied through word and pattern spotting, where the goal is to retrieve repeated words or writer specific patterns without full OCR. An efficient learning-free word spotting method based on local binary patterns was proposed for handwritten historical documents, demonstrating robustness to degradation and variations across multiple writers [7]. Bag of Visual Words (BoVW) and Bag of Words (BoW) models constitute a fundamental paradigm in image retrieval [19]. These approaches quantize local image descriptors into a shared codebook, representing each image as a histogram of visual words. Extending this to documents, Shekhar showed that recognition-free retrieval is possible by quantizing local descriptors and comparing fixed dimensional histograms [18]. A two stage system for word and pattern spotting in historical manuscripts combines BoVW for candidate generation with a spatial verification step based on the Longest Weighted Profile algorithm, enabling retrieval from a single query [8]. These methods are closely related, but they target repeated words or local patterns rather than page level manuscript identity. This global quantization can be challenging for fine grained scribal analysis: fragments with identical global word counts but entirely different geometric distributions in feature space. Optimal transport offers a more flexible alternative for comparing visual distributions. Earth Mover’s Distance (EMD) utilizes optimal transport to compare feature distributions [17], recently extended to match continuous GMMs [15]. We combine these threads: instead of assigning all pages to one global codebook, we learn a fragment-specific vocabulary and compare fragments through set to set matching.
@@
A closely related research area is writer identification [6], which seeks to determine the identity of the writer of a specific document based on handwriting out of a given set of writers for whom writing samples are supplied. Computer vision tools have been developed to assist paleographers directly in scribal hand identification, producing interpretable patch-level similarity maps that experts can interrogate alongside traditional paleographic evidence [12]. Existing methods are commonly divided into texture-based and grapheme based approaches. Texture-based methods [2, 3, 1] characterize handwriting using statistical properties of the written trace, such as slant, curvature, and texture, whereas grapheme based methods extract local structures and map them into a shared feature space. This latter can be formed using BoW, which relies on zero order statistics by counting assignments of local descriptors to the nearest visual words in a predefined vocabulary [9, 22] or by richer aggregation schemes such as Fisher vectors, which encode first and second order deviations of local descriptors from a probabilistic visual vocabulary [10, 13], and VLAD, which aggregates residuals between local descriptors and their nearest visual words [4, 5].
@@ 2.1 Dataset
Let $\mathcal{I}=\{I_{i}\}_{i=1}^{N}$ be a collection of manuscript fragment images partitioned into join clusters $\mathcal{C}=\{C_{j}\}_{j=1}^{M}$, where $C_{j}\subseteq\mathcal{I}$ denotes a set of fragment images originating from the same physical manuscript. Given a query fragment $I_{q}$, the retrieval task is to rank all other images so that members of $C_{j(q)}$, where $j(q)$ denotes the cluster index of $I_{q}$, appear at top ranks. We evaluate retrieval using Hit@$k$, mAP@$k$, mean reciprocal rank (MRR), and Macro-F1@1. The implemented system constructs two competing representations in a shared latent space: (i) an image adaptive Bag of Bags (BoB) representation, and (ii) a global codebook Bag of Words (BoW) baseline. The BoB pipeline proceeds in three stages: (1) character anchored patch extraction via connected component analysis; (2) sparse autoencoder encoding of each patch; and (3) per image $k$-means clustering to produce a local prototype vocabulary. Fragment similarity is then measured as a set to set distance between per image vocabularies.
@@ 2.3 Sparse Convolutional Autoencoder
where $\lambda=10^{-5}$ balances reconstruction fidelity and feature sparsity. For a fragment $I_{i}$, this produces a set of continuous embeddings $E_{i}=\{f_{\theta}(p):p\in\mathcal{P}_{i}^{*}\}\subseteq\mathbbm{R}^{128}$.
3 Methodology
3.1 Problem Formulation
@@ 3.3 Distance Formulation
where $\Pi(\boldsymbol{\pi},\boldsymbol{\rho})$ is the set of nonnegative transport plans with marginals $\boldsymbol{\pi}$ and $\boldsymbol{\rho}$, solved via the POT library [11]. A prototype accounting for 40% of a page’s components contributes proportionally more to the distance than one accounting for 2%. BoB-OT improves over $d_{\mathrm{Hung}}$ (+5.6 $p$ Hit@1), confirming that prototype mass is informative, though both assignment-based distances are ultimately outperformed by Chamfer in this highly-degraded domain.
@@ 3.4 Bag-of-Words (BoW) Baseline
We compare pages using Euclidean, cosine, $\chi^{2}$, and Hellinger distances. For Hellinger distance, histograms are first renormalized to unit $\ell_{1}$ mass. This baseline provides a meaningful shared-vocabulary comparator while discarding the page-specific prototype geometry retained by BoB.
@@ 3.5 Retrieval, Reranking and Complexity
For larger manuscript collections, where computing or storing a full $N{\times}N$ BoB-OT matrix becomes expensive—BoB-OT costs $\mathcal{O}(K^{3})$ per pair versus $\mathcal{O}(K_{g})$ for BoW—we propose a two-stage pipeline: BoW-Cosine retrieves the top-$M$ candidates efficiently, and BoB-OT reranks only that shortlist. We evaluate with M=30M{=}30, reducing online cost to $\mathcal{O}(M{\cdot}K^{3})$ per query independent of gallery size. See Fig. 2. Table 1 shows that the two-stage pipeline closely approximates exhaustive BoB-OT at Hit@1 and Hit@5; small differences at deeper ranks arise where BoW recall limits the candidate pool.
@@ 4.1 Retrieval Performance
Table 1 reports the full retrieval comparison across three families of baselines. Pooling baselines (MeanPool, MaxPool) aggregate all component embeddings into a single page vector without any clustering; MaxPool performs poorly (Hit@1 $\leq$ 0.31), confirming that element-wise maximum is not an informative aggregation for manuscript embeddings. MeanPool-Cosine is competitive at 0.545 Hit@1, but substantially below BoB, showing that a simple global average of component embeddings cannot substitute for structured page-level vocabulary matching. BoW-centroids constructs a global codebook over page-level BoB prototypes; its best variant (L2) achieves 0.545 Hit@1. BoW-Raw Patches is the standard Bag-of-Visual-Words formulation applied directly to raw component embeddings without any per-page clustering stage; its strongest variant ($\chi^{2}$) reaches 0.739 Hit@1 and 0.800 MRR, making it the most competitive baseline.
@@ 4.2 Distance-Space Separation
Table 2: Distance-separation statistics for intra-cluster ablation vs. inter-cluster pairs. Larger mean gap, KS statistic, and AUC separation indicate stronger distributional separation.
@@ 4.3 Ablation Studies
Table 3: Ablation on local vocabulary size $K$ for BoB-Hungarian-L2 ($d=128$, sparsity enabled)). (MF1 = MacroF1.)
@@
Table 4: Ablation on latent dimension $d$.
@@
Table 5 compares training with and without the $\ell_{1}$ latent penalty. Enabling sparsity improves Hit@1 (+0.011+0.011), Hit@5 (+0.023+0.023), mAP@5 (+0.020+0.020), and MRR (+0.002+0.002). MacroF1@1 also improves substantially under sparsity (+0.063+0.063). Overall, the sparse model is superior across all reported metrics; we retain the $\ell_{1}$ penalty in the default configuration.
@@ 5 Discussion
The evaluation is constrained to a manually annotated subset of 287 images forming 100 join clusters, reflecting the difficulty of obtaining verified, high-confidence ground-truth joins from domain experts. The complete Genizah corpus contains over 250,000 fragments, and the pipeline is designed with that full-corpus setting in mind. We expect the per-image vocabulary construction to remain advantageous at scale: adapting to each fragment’s local character distribution rather than a fixed global codebook becomes more valuable as stylistic diversity increases. Empirical validation on a larger annotated sample remains an important direction for future work, but awaits the accumulation of ground truth. We focus on BoW-family baselines sharing the same sparse encoder so as to isolate the effect of page-adaptive vocabulary construction; comparison against more recent learned retrieval models is another important direction for the future. Extending the pipeline to smaller fragments and more heavily degraded pages are natural directions for future work. On our benchmark test set, $K=32$, the approximate number of graphemes for those texts, yielded the strongest retrieval metrics, at non-negligible cost compared to the $K=20$ used in most of our experiments. It remains to be seen what the ideal number is for the complete corpus. It may in fact be best to form $K^{\prime}>K$ clusters and retain only the $K$ most populated prototypes for matching, decoupling vocabulary richness from matching cost.

### 10 Apr 2026 v2

@@ Abstract
A join is a set of manuscript fragments identified as originally emanating from the same manuscript. We study manuscript join retrieval: Given a query image of a fragment, retrieve other fragments originating from the same physical manuscript. We propose Bag of Bags (BoB), an image-level representation that replaces the global-level visual codebook of classical Bag of Words (BoW) with a fragment-specific vocabulary of local visual words. Our pipeline trains a sparse convolutional autoencoder on binarized fragment patches, encodes connected components from each page, clusters the resulting embeddings with per-image $k$-means, and compares images using set-to-set distances between their local vocabularies. Evaluated on fragments from the Cairo Genizah, the best BoB variant (viz. Chamfer) achieves Hit@1 of 0.78 and MRR of 0.84, compared to 0.74 and 0.80, respectively, for the strongest BoW baseline (BoW-RawPatches-$\chi^{2}$), a 6.1% relative improvement in top-1 accuracy. We furthermore study a mass-weighted BoB-OT variant that incorporates cluster population into prototype matching and present a formal approximation guarantee bounding its deviation from full component-level optimal transport. A two-stage pipeline using a BoW shortlist followed by BoB-OT reranking provides a practical compromise between retrieval strength and computational cost, supporting applicability to larger manuscript collections. The code and dataset are available at https://github.com/TAU-CH/midrash_bob.
@@ 1 Introduction
Automating join retrieval is a challenging computer vision task because many manuscript fragments are often severely damaged, incomplete, and stained. The difficulty arises from the fact that fragments originating from the same manuscript may share only subtle, handwriting-style-specific visual patterns, while fragments from different manuscripts can still appear globally similar in aspects such as background texture, aging, and degradation level.
Standard retrieval architectures frequently rely on Bag of Words (BoW) paradigms. BoW maps images to histograms over a shared global codebook. It summarizes all fragments with respect to the same dictionary. However, this global quantization is fundamentally limited for paleographic analysis, as it causes critical image-specific handwriting style information to be lost during quantization. Two images with identical codeword frequencies but different geometric character style distributions are assigned distance zero by BoW, yet they may have been written by entirely different hands. Conversely, fragments from the same manuscript may differ in lighting and layout, causing BoW to miss them. BoW’s pooling over a global codebook erases the image-specific handwriting style signal.
@@
  • We propose Bag of Bags (BoB), a fragment-specific representation for fragment retrieval that replaces a shared global visual-word codebook with per-fragment vocabularies of local visual words, constructed from sparse autoencoder embeddings of connected-component patches.
@@ 2 Related Work
Automated identification of joins was first explored in [21], who proposed a method for finding candidate matches between manuscript fragments. Their approach represents each image using local visual features and a bag of features representation, and then compares pairs of leaves using a learned similarity measure to determine whether they may belong to the same original work. However, this method requires computationally expensive pairwise comparisons during retrieval. Analogous fragment assembly problems arise in other historical document collections, including ancient papyri, where AI-based approaches have been surveyed for matching dispersed fragments via visual texture and learned similarity measures [20, 16].
@@
More broadly, historical document retrieval has been studied through word and pattern spotting, where the goal is to retrieve repeated words or writer-specific patterns without full OCR. An efficient learning-free word spotting method based on local binary patterns was proposed for handwritten historical documents, demonstrating robustness to degradation and variations across multiple writers [7]. Bag of Visual Words (BoVW) and Bag of Words (BoW) models constitute a fundamental paradigm in image retrieval [19]. These approaches quantize local image descriptors into a shared codebook, representing each image as a histogram of visual words. Extending this to documents, Shekhar showed that recognition-free retrieval is possible by quantizing local descriptors and comparing fixed-dimensional histograms [18]. A two-stage system for word and pattern spotting in historical manuscripts combines BoVW for candidate generation with a spatial verification step based on the Longest Weighted Profile algorithm, enabling retrieval from a single query [8]. These methods are closely related, but they target repeated words or local patterns rather than page-level manuscript identity. This global quantization can be challenging for fine-grained scribal analysis: fragments with identical global word counts but entirely different geometric distributions in feature space receive identical representations. Optimal transport offers a more flexible alternative for comparing visual distributions. Earth Mover’s Distance (EMD) utilizes optimal transport to compare feature distributions [17], recently extended to match continuous GMMs [15]. We combine these threads: instead of assigning all pages to one global codebook, we learn a fragment-specific vocabulary and compare fragments through set-to-set matching.
@@
A closely related research area is writer identification [6], which seeks to determine the identity of the writer of a specific document based on handwriting out of a given set of writers for whom writing samples are supplied. Computer vision tools have been developed to assist paleographers directly in scribal hand identification, producing interpretable patch-level similarity maps that experts can interrogate alongside traditional paleographic evidence [12]. Existing methods are commonly divided into texture-based and grapheme-based approaches. Texture-based methods [2, 3, 1] characterize handwriting using statistical properties of the written trace, such as slant, curvature, and texture, whereas grapheme-based methods extract local structures and map them into a shared feature space. The latter can be formed using BoW, which relies on zero-order statistics by counting assignments of local descriptors to the nearest visual words in a predefined vocabulary [9, 22] or by richer aggregation schemes such as Fisher vectors, which encode first and second-order deviations of local descriptors from a probabilistic visual vocabulary [10, 13], and VLAD, which aggregates residuals between local descriptors and their nearest visual words [4, 5].
3 Methodology
3.1 Problem Formulation
Let $\mathcal{I}=\{I_{i}\}_{i=1}^{N}$ be a collection of manuscript fragment images partitioned into join clusters $\mathcal{C}=\{C_{j}\}_{j=1}^{M}$, where $C_{j}\subseteq\mathcal{I}$ denotes a set of fragment images originating from the same physical manuscript. Given a query fragment $I_{q}$, the retrieval task is to rank all other images so that members of $C_{j(q)}$, where $j(q)$ denotes the cluster index of $I_{q}$, appear at top ranks. We evaluate retrieval using Hit@$k$, mAP@$k$, mean reciprocal rank (MRR), and Macro-F1@1. The implemented system constructs two competing representations in a shared latent space: (i) an image-adaptive Bag of Bags (BoB) representation, and (ii) a global codebook Bag of Words (BoW) baseline. The BoB pipeline proceeds in three stages: (1) character-anchored patch extraction via connected component analysis; (2) sparse autoencoder encoding of each patch; and (3) per-image $k$-means clustering to produce a local prototype vocabulary. Fragment similarity is then measured as a set-to-set distance between per-image vocabularies.
@@ 3.4 Sparse Convolutional Autoencoder
where $\lambda=10^{-5}$ balances reconstruction fidelity and feature sparsity. For a fragment $I_{i}$, this produces a set of continuous embeddings $E_{i}=\{f_{\theta}(p):p\in\mathcal{P}_{i}\}\subseteq\mathbbm{R}^{128}$.
@@ 3.6 Distance Formulation
where $\Pi(\boldsymbol{\pi},\boldsymbol{\rho})$ is the set of non-negative transport plans with marginals $\boldsymbol{\pi}$ and $\boldsymbol{\rho}$, solved via the POT library [11]. A prototype accounting for 40% of a page’s components contributes proportionally more to the distance than one accounting for 2%. BoB-OT improves over $d_{\mathrm{Hung}}$ (+5.6 $p$ Hit@1), confirming that prototype mass is informative, though both assignment-based distances are ultimately outperformed by Chamfer in this highly-degraded domain.
@@ 3.7 Bag-of-Words (BoW) Baseline
We compare pages via distances between their final page-level BoW representations: Euclidean, cosine, $\chi^{2}$, and Hellinger. For Hellinger distance, histograms are first renormalized to unit $\ell_{1}$ mass. This baseline provides a meaningful shared-vocabulary comparator while discarding the page-specific prototype geometry retained by BoB.
@@ 3.8 Retrieval, Reranking and Complexity
For larger manuscript collections, where computing or storing a full $N{\times}N$ BoB-OT matrix becomes expensive—BoB-OT costs $\mathcal{O}(K^{3})$ per pair versus $\mathcal{O}(K_{g})$ for BoW—we propose a two-stage pipeline: BoW-Cosine retrieves the top-$M$ candidates efficiently, and BoB-OT reranks only that shortlist. We evaluate with M=30M{=}30, reducing online cost to $\mathcal{O}(M{\cdot}K^{3})$ per query independent of gallery size. See Fig. 2. Table 1 shows that the two-stage pipeline matches or improves exhaustive BoB-OT across all reported ranking metrics on this benchmark.
@@ 4.1 Retrieval Performance
Table 1 reports the full retrieval comparison across three families of baselines. Pooling baselines (MeanPool, MaxPool) aggregate all component embeddings into a single page vector without any clustering; MaxPool performs poorly (Hit@1 $\leq$ 0.31), confirming that the element-wise maximum is not an informative aggregation for manuscript embeddings. MeanPool-Cosine is competitive at 0.545 Hit@1, but substantially below BoB, showing that a simple global average of component embeddings cannot substitute for structured page-level vocabulary matching. BoW-centroids constructs a global codebook over page-level BoB prototypes; its best variant (L2) achieves 0.545 Hit@1. BoW-RawPatches is the standard Bag-of-Visual-Words formulation applied directly to raw component embeddings without any per-page clustering stage; its strongest variant ($\chi^{2}$) reaches 0.739 Hit@1 and 0.800 MRR, making it the most competitive baseline.
@@ 4.2 Distance-Space Separation
Table 2: Distance-separation statistics for intra-cluster vs. inter-cluster pairs. Larger mean gap, KS statistic, and AUC separation indicate stronger distributional separation.
@@ 4.3 Ablation Studies
Table 3: Ablation on local vocabulary size $K$ for BoB-Hungarian-L2 ($d=128$, sparsity enabled). (MF1 = MacroF1)
@@
Table 4: Ablation on latent dimension $d$ (K=32, sparsity enabled).
@@
Table 5 compares training with and without the $\ell_{1}$ latent penalty. Enabling sparsity improves Hit@1 (+0.012+0.012), mAP@1 (+0.012+0.012), Hit@5 (+0.023+0.023), mAP@5 (+0.017+0.017), and MRR (+0.008+0.008). MacroF1@1 also improves substantially under sparsity (+0.052+0.052). Overall, the sparse model is superior across all reported metrics; we retain the $\ell_{1}$ penalty in the default configuration.
@@ 5 Discussion
The evaluation is constrained to a manually annotated subset of 287 images forming 100 join clusters, reflecting the difficulty of obtaining verified, high-confidence ground-truth joins from domain experts. The complete Genizah corpus contains over 250,000 fragments, and the pipeline is designed with that full-corpus setting in mind. We expect the per-image vocabulary construction to remain advantageous at scale: adapting to each fragment’s local character distribution rather than a fixed global codebook becomes more valuable as stylistic diversity increases. Empirical validation on a larger annotated sample remains an important direction for future work, but awaits the accumulation of ground truth. We focus on BoW-family baselines sharing the same sparse encoder so as to isolate the effect of page-adaptive vocabulary construction; comparison against more recent learned retrieval models is another important direction for the future. Extending the pipeline to smaller fragments and more heavily degraded pages is a natural direction for future work. On our benchmark test set, $K=32$, the approximate number of graphemes for those texts, yielded the strongest retrieval metrics, at non-negligible cost compared to the $K=20$ used in most of our experiments. It remains to be seen what the ideal number is for the complete corpus. It may in fact be best to form $K^{\prime}>K$ clusters and retain only the $K$ most populated prototypes for matching, decoupling vocabulary richness from matching cost.